you want to serve 210 people each a 6 fluid ounce portion of orange juice. orange juice concentrate is purchased by the case for $22.80. each case contains 12 cans, and each can makes 3/4 gallon. how much will the juice cost for the 210 people?
Answer by LinnW(1048)

Let's first determine the number of ounces required.
( 210 people * 6 oz ) = 1260 oz
Now determine how many cases are required to produce 1260 oz.
Each case makes 3/4 gallon. Since 1 gallon = 128 oz ,
3/4 gallon = (3/4)*128 = 3*128/4 = 3*32 = 96 oz.
We now know that each case produces 96 oz of juice.
If we divide 1260 by 96 this gives the number of cases required.
1260/96 = 13 1/8 cases .
Since we buy juice in cases, we will need 14 cases.
( 14 cases )*( $22.80 per case ) = $319.20